OpenStudy (abmon98):

(2sintheta*costheta)+sintheta=0

OpenStudy (abmon98):

solve equation for the domain

OpenStudy (abmon98):

\[2\sin \theta \cos \theta+\sin \theta=0\]

OpenStudy (anonymous):

\[\sin \theta \left( 2\cos \theta+1 \right)=0\] \[Either~\sin \theta=0=\sin k \pi,k \in I(integers)\] \[\theta=k \pi \] \[\cos \theta=\frac{ -1 }{ 2 }=-\cos \frac{ \pi }{3 }=\cos \left( 2k \pi+\pi-\frac{ \pi }{ 3 } \right),\cos \left( 2k \pi+\pi+\frac{ \pi }{ 3 } \right)\] \[\theta=\left( 2k+\frac{ 2 }{ 3 } \right)\pi,\left( 2k+\frac{ 4 }{3 } \right)\pi,k \in I\]